Equipment:- OSN 3500
This NE is connected with other NE with SL16 cards.,
There were B1_SD,B2_SD and B3_SD alarms on 13-SL16 card of OSN 3500.
After checked the received optical power at SL16 card and found that its received power was very low due to high fiber loss. When received power came to specified range then these alarms have cleared.